A bucket full of hot water is kept in a room and it cools from 75°C to 70°C in t1 minutes from 70°C to 65°C in t2 minutes and from 65°C to 60°C in t3 minutes; then

(a) t1 - t2 = t3
(b) t1 < t2 < t3
(c) t1 > t2 > t3
(d) t1 < t2 > t3

Grade:Upto college level

3 Answers

Priyanka motiani
7 years ago
T1-T2/time = k ( T1+T2/2 - T(surronding)) now put the values of temp and time. Keep time and tsurr as it is in the form of variable only we have 3 different time so we get 3 eq. .5/t1 = k (135/2- Tsu) ,5/t2 = k (125/2-Tsu), .5/t3 =k (115/2 -Tsu) divide two eq at a time the you get a idea about t1, t2 and t3 so you get option b ... hope it helps
ankit singh
3 years ago
A bucket full of hot water is kept in a room and it cools from 75°C to 70°C in T1 minutes, from 70°C to 65°C in T_2 minutes and from 65°C to 60°C in T_3 minutes.

 
ankit singh
3 years ago
thank u regrds ankit singh
this questionis so so easy byte way u firstly u just read the
question carefully